Prof. Theodore Gamelin Mathematics Department UCLA 405 Hilgard Avenue Los Angeles, CA 90095-1555 USA

Composition Operators on Uniform Algebras

Abstract. The talk will focus on composition operators on uniform algebras and the existence of attracting cycles for the corresponding self-map of the spectrum of the uniform algebra. The pseudohyperbolic metric on the spectrum and hyperbolically bounded sets will play a decisive role.